# Depression Detection on Social Media with Large Language Models

### v1

@@ Abstract
Depression harms. However, due to a lack of mental health awareness and fear of stigma, many patients do not actively seek diagnosis and treatment, leading to detrimental outcomes. Depression detection aims to determine whether an individual suffers from depression by analyzing their history of posts on social media, which can significantly aid in early detection and intervention. It mainly faces two key challenges: 1) it requires professional medical knowledge, and 2) it necessitates both high accuracy and explainability. To address it, we propose a novel depression detection system called DORIS, combining medical knowledge and the recent advances in large language models (LLMs). Specifically, to tackle the first challenge, we proposed an LLM-based solution to first annotate whether high-risk texts meet medical diagnostic criteria. Further, we retrieve texts with high emotional intensity and summarize critical information from the historical mood records of users, so-called mood courses. To tackle the second challenge, we combine LLM and traditional classifiers to integrate medical knowledge-guided features, for which the model can also explain its prediction results, achieving both high accuracy and explainability. Extensive experimental results on benchmarking datasets show that, compared to the current best baseline, our approach improves by 0.036 in AUPRC, which can be considered significant, demonstrating the effectiveness of our approach and its high value as an NLP application.
@@ 1 Introduction
Depression profoundly affects humanity, with WHO estimates indicating that 5% of adults suffer from it111www.who.int/news-room/fact-sheets/detail/depression, significantly contributing to the global suicide rate222www.ncbi.nlm.nih.gov/pmc/articles/PMC6165520/. Given the harmful nature of depression, timely diagnosis and intervention are necessary. However, traditional hospital-based approaches for diagnosing depression face several issues. Firstly, patients often avoid evaluations due to stigma or not recognizing their need for help rusch2005mental . What’s more, self-reported diagnoses can be unreliable due to intentional concealment bril2017qualitative . Additionally, the high cost of hospital evaluations places a burden on patients khan2016economic . For these reasons, many individuals with depression remain undiagnosed and untreated, with over 75% of those in low- and middle-income countries receiving no treatment at all333press.un.org/en/2021/sgsm20951.doc.htm.
Depression detection on social media identifies potential depression through users’ post histories orabi2018deep ; shen2017depression , as shown in Figure 1, which is one typical beneficial application of NLP techniques. That is, it leverages public posts from online social networks for broad detection coverage, benefits from more genuine expressions than those in clinical settings lachmar2017mydepressionlookslike , and reduces economic costs compared to professional diagnoses.
Figure 1: Illustration of depression detection on social media. It aims to determine whether a user has depression based on their history of posts on social media.
Many works have delved into depression detection based on texts from social media. eichstaedt2018facebook first identified the possibility of predicting medically diagnosed depression through expressions in social media posts. Recent works lin2020sensemood ; orabi2018deep ; yang2022mental ; chiong2021combining utilize deep learning methods to determine whether users suffer from depression based on their historical posts. However, they are not able to incorporate medical knowledge on depression, leaving significant room for improvement in both accuracy and interpretability. With the advent of LLMs, some studies have employed them for mental health analysis, achieving good explainability yang2023towards . However, these methods fall short in accuracy arcan2024assessment , and are highly sensitive to small prompt variations hua2024large .
To summarize, previous work has not resolved two significant challenges in depression detection as follows.
  • First, depression detection necessitates reference to professional medical knowledge. Extensive research in medicine has led to widely used diagnostic methods and standards, including depression scales and mood course analysis. Integrating these explicit standards into automated systems presents a challenge, moving beyond the common use of black-box approaches like embedding models. It requires a nuanced analysis of posts and targeted modeling strategies.
  • Second, depression detection demands both high accuracy and high explainability. As a serious and safety-critical task, it requires trustworthy methods with high performance in both two aspects. While traditional non-LLM methods often achieve higher accuracy, they typically lack explainability. In contrast, LLMs naturally explain their judgments but tend to be less accurate. Thus, designing an approach that combines their strengths, ensuring accuracy and explainability, is essential yet very challenging.
To address these challenges, in this work, we propose a depression-detection system named DORIS (short for DiagnOstic CRiteria-Guided Mood HIStory-Aware). To tackle the first challenge, we incorporate the widely used depression scale DSM-V. Leveraging the strong text understanding capability, we adopt an LLM to assess whether each user’s high-risk texts reflect symptoms of depression, thereby constructing expert features for each user. Moreover, to model users’ mood courses, we use a text embedding model to filter posts with high emotional intensity and summarize these texts into users’ mood courses using an LLM. These high-emotion texts and mood course descriptions are then vectorized to generate representations of mood course. To tackle the second challenge, we combine features from different spaces to produce our method’s final judgment by using the GBT classifier, achieving high accuracy. Additionally, using the annotations of high-risk texts and descriptions of mood history generated during the model’s operation, we create a comprehensive explanation of the system’s judgment, achieving high explainability.
To validate the effectiveness of our method, we conduct extensive experiments, and our method shows an improvement of 0.036 over the current best baseline on the AUPRC metric. The ablation study demonstrates the effectiveness of each component of our method. The case study illustrates that our method can generate explainable diagnostic reports, achieving high explainability.
Our contribution can be summarized as follows:
(1) We are among the first to combine professional medical knowledge and advanced large language models in depression detection.
(2) We propose the depression detection system DORIS, aiming to integrate professional medical knowledge with advanced NLP techniques to generate accurate and interpretable judgments.
(3) We conduct extensive experiments demonstrating the accuracy of DORIS, the effectiveness of each module’s design, and the high quality of the model’s output explanations.
The organization of our paper is as follows. In Section 2, we review related work. Then, in Section 3, we clearly describe the problem definition and our method design. Afterward, in Section 4, we present our experimental results. Finally, in Section 5, we summarize our work and propose future directions.
2 Related Works
2.1 Depression Detection on Social Media
Compared to traditional medical diagnostic methods for depression conducted in hospitals, depression detection on social media has the advantages of lower concealment potential, wider coverage, and lower cost malhotra2022deep . Early research works first extract text features, then apply traditional machine learning methods for classification. Common feature extraction methods in depression detection research include LIWC tausczik2010psychological , TF-IDF, LDA blei2003latent , etc.; classifiers include SVM tadesse2019detection , Logistic Regression coppersmith2014quantifying , Random Forest cacheda2019early , among others. In recent years, with the development of deep learning technologies, deep learning techniques have been used in depression detection, such as CNN cohan2018smhd , RNN ive2018hierarchical , PLM owen2020towards ; ji2023domain , GCN naseem2022early . These efforts have improved the accuracy of depression detection, advancing the field’s research.
Some works on depression detection focus solely on analyzing individual texts to determine the presence of depressive symptoms yang2022mental ; chiong2021combining , the type of symptoms displayed zhang2023phq , and the level of depression zhang2023sentiment . However, analyzing a user’s post history is more informative, as it is common for individuals not suffering from depression to also occasionally post texts that exhibit depressive symptoms. Recent approaches to depression detection based on a user’s post history, such as those by eichstaedt2018facebook , orabi2018deep , lin2020sensemood , chiong2021combining , and yang2022mental , lack comprehensive integration of medical knowledge, limiting their accuracy and interpretability. Our research stands out as one of the initial efforts to thoroughly apply medical insights, leading to improved accuracy. Additionally, we are pioneers in employing LLMs for crafting systematic explanations of predictions, thereby increasing interpretability.
2.2 Large Language Models for Mental Health Analysis
Large language models (LLMs) have shown great potential in clinical applications due to its abundant prior knowledge and strong language generalization ability singhal2023large . Recent works have introduced LLMs into mental health analysis. arcan2024assessment presents a solid evaluation on the performance of Llama-2 touvron2023llama and ChatGPT [^4] https://openai.com/blog/chatgpt in mental health detection tasks, unveiling prospects along with challenges of LLM-based methods in mental health analysis. yang2023towards further discover the applications of LLMs in both mental health detection tasks and reasoning tasks, which highlight LLM’s excellent interpretability. yang2023mentalllama propose the first open-source LLM series for mental health analysis based on Llama-2 and fine-tuning techniques, and greatly enhances the accuracy and explanation quality compared with general-purpose LLMs. However, the challenge of LLM-based methods being weaker in terms of prediction accuracy compared to embedding models, especially when targeting specific downstream tasks, remains unresolved arcan2024assessment .
To our knowledge, our system is the first to be specifically designed for depression detection in the context of LLM for mental health analysis. Furthermore, we have facilitated a collaboration between LLMs and embedding models, attaining both high accuracy and interpretability, addressing deficiencies in previous designs.
@@ 3.1 Problem Formulation
In this study, we aim to assess the risk of depression in individuals based on their post history. For a user $u$, we can collect their historical posts, denoted as $P=\{p_{1},p_{2},...,p_{n}\}$. Each of the posts has a corresponding timestamp, denoted as $t_{1},t_{2},...,t_{n}$. Our task is formalized as determining the depression label $y$ for user $u$, given their history of posts $P$, where $y\in{0,1}$.
3.2 Model Design
In this part, we describe our proposed DORIS in detail. The architecture of DORIS is illustrated in Figure 2.
3.2.1 Diagnostic Criteria Feature Extraction
@@ I. Thoughts of suicide
Annotation with LLM
In the medical domain, there are well-established diagnostic frameworks for depression, yet its full potential has not been tapped into for depression detection on social media. Our system seeks to incorporate these diagnostic criteria, enabling decisions supported by medical knowledge, thus improving decision accuracy and credibility. Specifically, we build user features based on medical diagnostic criteria,
We utilize the DSM-5 regier2013dsm , a widely recognized tool for mental disorder assessment and diagnosis. It offers a list of clinical criteria for depression, detailed in Table 1.
LLMs exhibit a remarkable capacity for semantic understanding achiam2023gpt , and research has demonstrated their potential to replace human annotators in certain tasks ziems2023can . Here, we employ LLMs to annotate texts, specifically to identify if and which self-expressed symptoms of depression are present in posts. The prompt is structured as follows:
Assuming you are a psychiatrist specializing in depression. Given ‘text’, please determine if this message includes any of the following states of the author:
A. Depressive mood B. Loss of interest/pleasure … I. Thoughts of suicide.
If present, answer in the format of enclosed letters separated by commas, for example, (A, B, C). If none are present, respond with None.
@@
Efficient Annotation
The use of LLMs incurs substantial financial and energy costs. Since the vast majority of posts on social media platforms are unrelated to depression, annotating all posts would be exceedingly wasteful. To address this issue, we design an efficient annotation approach that first filters for high-risk texts and then annotates only those texts. Our experiments will demonstrate that annotating just the selected high-risk texts achieves a performance comparable to annotating all texts, while significantly reducing the number of requests to the LLM. Specifically, for each symptom, we designed a symptom template containing potential first-person textual expressions of that symptom. For example, for the symptom B. Loss of interest/pleasure, the template $T^{DC}_{B}$ is:
@@
3.2.2 Mood Course Representation Construction
Mood course, defined as the temporal pattern and progression of emotional states cochran2016data , is critical in diagnosing clinical depression costello2002development . It delineates the onset, duration, and recurrence of mood episodes, providing insights into the disorder’s nature and trajectory frias2017longitudinal . Accurately modeling mood course is pivotal for distinguishing depressive disorders from transient mood fluctuations, facilitating early detection and appropriate intervention rudolph2006mood .
Former works on depression detection has largely overlooked the mood course, focusing instead on static mood snapshots. Our study bridges this gap by explicitly modeling mood course and integrating it into our classification system. Next, we detail our approach.
Posts Filtering
Not all posts are emotionally charged. We begin by filtering posts with a high emotional content. Following oatley1987towards , we categorize emotions into five main types: 1) anger, 2) disgust, 3) anxiety, 4) happiness, and 5) sadness. For each of these emotional categories, we establish a template of emotional expressions. For instance, the template for sadness, $T_{5}^{E}$, is defined as:
@@ 3.2.2 Mood Course Representation Construction
For each user $u$, we intersect their historical post set $P_{u}$ with the high emotional content set $S$ to obtain $P_{u}^{E}$. Based on this subset of emotionally expressive posts, we use an LLM to synthesize a description of user $u$’s mood course, $T^{MC}$, using the following prompt:
"Assuming you are a psychiatrist specializing in depression, please summarize the mood course over this period based on the blogger’s self-expressions:
Time: $t_{1}$, Post: $p_{1}$, Time: $t_{2}$, Post: $p_{2}$, …"
@@
The mood course representation $F_{u}^{MC}$ for user $u$, is calculated as:
$F_{u}^{MC}=\frac{1}{\left|P_{u}^{E}\right|+1}\left(\sum_{p\in P_{u}^{E}}H_{p}+H^{MC}\right),$ (8)
where $\left|P_{u}^{E}\right|$ is the total number of posts in $P_{u}^{E}$.
3.2.3 Post History Representation Construction
In the sections above, we construct the diagnostic criteria feature and mood course representation from users’ historical posts through filtering and labeling. While these filtering steps emphasize aspects crucial for medical diagnosis of depression, they may also result in information loss. To address this, we constructed a representation of the user’s post history as follows:
@@ 3.2.3 Post History Representation Construction
3.2.4 Training and Predicting
In this section, we describe our training and predicting methodology. First, we integrate the various features. $F^{MC}$ and $F^{PH}$, which shares the same space, are directly summed to avoid increasing dimension and exacerbating the risk of overfitting. Conversely, $F^{DC}$ resides in a distinct space, and thus we concatenate this feature with the sum of the first two parts. The final feature vector $F$ is obtained as follows:
@@ 3.2.4 Training and Predicting
We employ a Gradient Boosting Trees (GBT) approach for classification. This method involves constructing an ensemble of decision trees in a sequential manner, where each subsequent tree aims to correct the errors of its predecessor. It excels in automatically performing feature interactions by selecting optimal split criteria within its decision trees, thus effectively fusing the components of $F$. The final prediction, denoted as $y$, is a binary classification result derived from the ensemble model. The process is formalized as follows:
@@
3.2.5 Design for Explainability
Depression detection is a critical and safety-sensitive task, demanding both accuracy and explainability zhang2022natural . Existing work has revealed the potential of large models in interpretable mental health analysis yang2023towards ; yang2023mentalllama . However, directly analyzing raw texts with LLMs can lead to accuracy discrepancies with domain-specific models and highly prompt-sensitive predictions hua2024large . In our approach, to ensure accuracy and stability, we do not rely on LLMs for direct judgment but instead utilize traditional classifiers. However, we also leverage the capabilities of LLMs to provide reasonable explanations for our model’s decisions.
In our system, we employ LLMs to annotate symptoms of depression manifested in texts posted on social media platforms by users. Additionally, we generate descriptions of users’ mood courses. These outputs serve not merely as intermediate variables in the system’s operation but also as part of the final output, enhancing users’ understanding of the system’s results. Furthermore, we use the LLM to generate a new output, $T^{E}xp$, offering explanations for the model’s classification results. The prompt used is as follows:
"Assuming you are a psychiatrist specializing in depression. Here is a user’s mood course: $T^{MC}$; below are posts from this user displaying symptoms of depression and the types of symptoms exhibited: …; this user has been determined by an automated depression detection system to be depressed/normal. Please consider the user’s mood course and posts to generate an explanation for this judgment."
This explanatory output, $T^{Exp}$, along with the system’s annotations of depressive symptoms in tweets, user mood course descriptions, and the system’s classification results, constitute the final output of our system.
4 Experiments
4.1 Experimental Setup
4.1.1 Implementation Details

### v2

@@ Abstract
Limited access to mental healthcare resources hinders timely depression diagnosis, leading to detrimental outcomes. Social media platforms present a valuable data source for early detection, yet this task faces two significant challenges: 1) the need for medical knowledge to distinguish clinical depression from transient mood changes, and 2) the dual requirement for high accuracy and model explainability. To address this, we propose DORIS, a framework that leverages Large Language Models (LLMs). To integrate medical knowledge, DORIS utilizes LLMs to annotate user texts against established medical diagnostic criteria and to summarize historical posts into temporal mood courses. These medically-informed features are then used to train an accurate Gradient Boosting Tree (GBT) classifier. Explainability is achieved by generating justifications for predictions based on the LLM-derived symptom annotations and mood course analyses. Extensive experimental results validate the effectiveness as well as interpretability of our method, highlighting its potential as a supportive clinical tool.
[…]
Depression Detection on Social Media with Large Language Models
Xiaochong Lan1thanks: Equal contribution. Zhiguang Han211footnotemark: 1 Yiming Cheng3 Li Sheng1 Jie Feng1 Chen Gao1thanks: Corresponding author. Yong Li122footnotemark: 2 1Department of Electronic Engineering, BNRist, Tsinghua University 2College of Computing and Data Science, Nanyang Technological University 3Department of Computer Science, The University of Chicago lanxc22@mails.tsinghua.edu.cn {chgao96, liyong07}@tsinghua.edu.cn
@@ 1 Introduction
Depression is a major global health challenge, yet significant barriers like cost, stigma, and lack of infrastructure prevent many from receiving timely care, creating a critical treatment gap Organization ; Khan et al. (2016); Rüsch et al. (2005). To help bridge this gap, analyzing public social media data offers a scalable approach for early detection. By examining patterns in users’ posts, data-driven systems can provide clinicians with timely signals of potential distress, augmenting traditional diagnostic methods Orabi et al. (2018); Shen et al. (2017); Sarkar et al. (2022).
However, building a clinically valid and trustworthy system from this data presents significant hurdles. Many existing deep learning methods (Lin et al., 2020; Orabi et al., 2018; Sarkar et al., 2022) lack a systematic integration of medical knowledge, struggling to distinguish the complex patterns of clinical depression from transient sadness Zhang et al. (2023a). Furthermore, they often function as “black boxes”, failing to provide the transparent, evidence-based reasoning that clinicians require for validation and trust. While Large Language Models (LLMs) offer powerful interpretation capabilities Yang et al. (2023a), their direct use for classification can lead to lower accuracy and reliability issues Arcan et al. (2024); Hua et al. (2024). This reveals a critical need for a framework that synergistically combines clinical knowledge, robust prediction, and explainability.
Figure 1: Depression detection on social media analyzes users’ social media post history to identify potential depression, offering insights to support mental health professionals in depression assessment.
To address these challenges, we propose DORIS (short for Dia-gnOstic CRiteria-Guided Mood HIStory-Aware). For the first challenge, we leverage LLMs’ strong language understanding capabilities to recognize complex expressions of depression symptoms based on DSM-5 criteria. We also systematically model users’ mood courses by filtering emotionally intensive posts and using LLMs to analyze their temporal patterns, enabling holistic assessment of users’ emotional fluctuations. For the second challenge, we design a hybrid approach that combines LLMs for explainable symptom and mood course analysis with GBT classifier for a robust final prediction. This design achieves high accuracy through the classifier while maintaining explainability in two ways: through concrete evidence derived from medical knowledge (annotated symptoms and mood courses), and through LLM-generated explanations that connect this evidence to depression assessment.
Our method shows significant improvements in experimental evaluation, achieving a 0.0361 gain in AUPRC over the current best baseline. Ablation studies validate each component’s contribution, and case studies demonstrate the system’s ability to generate explainable assessments backed by concrete evidence.
Our contributions can be summarized as follows:
  • We propose a novel hybrid depression detection framework that synergistically combines a robust classifier for accurate prediction with an LLM for explainable, clinically-grounded feature generation, addressing the critical dual need for accuracy and interpretability.
  • We introduce a methodology to systematically operationalize medical knowledge by using LLMs for both fine-grained DSM-5 symptom annotation and longitudinal mood course analysis, creating features that align closely with clinical diagnostic practices.
  • We demonstrate through extensive experiments on real-world data that our approach not only significantly improves detection performance but also provides clinically interpretable explanations grounded in medical knowledge, demonstrating a viable pathway for creating practical AI-driven tools that support clinicians and improve mental healthcare access in low-resource settings.
@@ 2.1 Problem Formulation
This study addresses the task of depression detection based on users’ social media posts. Users often use posts to document events around them and express their feelings, making these candid expressions valuable for gauging the likelihood of depression. Given a user $u$, let $P=\{p_{1},p_{2},...,p_{n}\}$ denote their historical posts with the corresponding timestamps $\{t_{1},t_{2},...,t_{n}\}$. Our objective is to learn a mapping function $f:P\rightarrow y$ that predicts the depression status $y\in\{0,1\}$ of user $u$, where $y=1$ indicates the presence of depression and $y=0$ indicates its absence.
In the following part, we describe our proposed DORIS in detail. The architecture of DORIS is illustrated in Figure 2.
2.2 Diagnostic Criteria Feature Extraction
2.2.1 Annotation with LLM
Clinical practice offers well-established diagnostic frameworks for depression that have proven highly effective in professional assessment. To bring this medical expertise to social media-based detection, we incorporate standardized diagnostic criteria into our system’s design. Specifically, we utilize DSM-5 (Regier et al., 2013), a widely recognized diagnostic tool in psychiatry that provides comprehensive criteria for mental disorder assessment and diagnosis. The core criteria in DSM-5, which guide mental health professionals in their diagnostic practice, are detailed in Table 1.
@@ I. Thoughts of suicide
Prior attempts to integrate clinical standards via DSM-5 criteria or its simplified screening tool PHQ-9 Yadav et al. (2020); Nguyen et al. (2022); Zhang et al. (2023a) rely on simple classification models that struggle to comprehend complex symptom expressions in social media texts. The emergence of large language models (LLMs) offers a promising solution to this challenge. LLMs exhibit a remarkable capacity for semantic understanding (Achiam et al., 2023), and research has demonstrated their potential to replace human annotators in certain tasks (Ziems et al., 2023). Here, we employ LLMs to accurately annotate texts, specifically to identify if and which self-expressed symptoms of depression are present in posts. The prompt is described in Appendix F.
@@
2.2.2 Efficient Annotation.
The use of LLMs incurs substantial financial and energy costs. Since the vast majority of posts on social media platforms are unrelated to depression, annotating all posts would be exceedingly wasteful. To address this issue, we design an efficient annotation approach that first filters for high-risk texts through carefully validated symptom templates (detailed in Appendix G) and then annotates only those texts with LLMs. We developed symptom templates through a four-stage process. First, psychiatrists created initial templates from DSM-5 diagnostic criteria. Second, we refined them using a corpus of 500 social media posts independently annotated by four clinicians, evaluating inter-annotator agreement (Cohen’s kappa) to improve them. Third, experts iteratively reviewed the templates for clinical accuracy and coverage until reaching consensus. Finally, experimental validation confirmed that filtering with our templates achieves comparable performance to full LLM annotation at a significantly lower computational cost.
@@
2.3 Mood Course Representation Construction
The mood course, also termed mood trajectory in clinical literature, represents the temporal pattern and progression of emotional states (Cochran et al., 2016) and is critical in diagnosing clinical depression (Costello et al., 2002). It delineates the onset, duration, and recurrence of mood episodes, providing insights into the disorder’s nature and trajectory (Frías et al., 2017). Accurately modeling mood course is pivotal for distinguishing depressive disorders from transient mood fluctuations, facilitating early detection and appropriate intervention (Rudolph et al., 2006). However, former works on depression detection have largely overlooked the mood course, focusing instead on static mood snapshots. Our study bridges this gap by explicitly modeling mood course and integrating it into our classification system. Next, we detail our approach.
2.3.1 Posts Filtering
Not all posts are emotionally charged. We begin by filtering posts with a high emotional content. Following Oatley and Johnson-Laird (1987), we categorize emotions into five main types: 1) anger, 2) disgust, 3) anxiety, 4) happiness, and 5) sadness. For each of these emotional categories, with the help of professional psychologists, we establish a template of emotional expressions (detailed in Appendix H) using the same systematic four-stage process described for depression symptom templates. For instance, the template for sadness, $T_{5}^{E}$, is defined as:
@@ 2.3.2 Representation Construction
For each user $u$, we intersect their historical post set $P_{u}$ with the high emotional content set $S$ to obtain $P_{u}^{E}$. Based on this subset of emotionally expressive posts, we use an LLM to synthesize a description of user $u$’s mood course, $T^{MC}$, the prompt is described in Appendix F.
@@
The user $u$’s mood course representation $F_{u}^{MC}$ is computed as:
$F_{u}^{MC}=\alpha H^{MC}+\beta\frac{1}{\left|P_{u}^{E}\right|}\sum_{p\in P_{u}^{E}}H_{p},$ (8)
where $\left|P_{u}^{E}\right|$ is the total number of posts in $P_{u}^{E}$, and $\alpha$, $\beta$ are hyperparameters. Here we get $F_{u}^{MC}$ as a comprehensive representation of the types and evolution of user historical moods.
2.4 Post History Representation Construction
In the sections above, we construct the diagnostic criteria feature and mood course representation from users’ historical posts through filtering and labeling. While these filtering steps emphasize aspects crucial for medical diagnosis of depression, they may also result in information loss. To address this, we construct a representation of the user’s post history as follows:
@@ 2.4 Post History Representation Construction
2.5 Training and Predicting
In this section, we describe our training and predicting methodology. First, we integrate the various features. $F^{MC}$ and $F^{PH}$, which shares the same space, are directly summed to avoid increasing the dimension and exacerbating the risk of overfitting. Conversely, $F^{DC}$ resides in a distinct space, and thus we concatenate this feature with the sum of the first two parts. The final feature vector $F$ is obtained as:
@@ 2.5 Training and Predicting
We employ Gradient Boosting Trees (GBT) for classification. GBT is an ensemble learning method that constructs a sequence of decision trees, where each subsequent tree aims to correct the errors of its predecessor. It excels in automatically performing feature interactions by selecting optimal split criteria within its decision trees, thus effectively fusing the components of $F$. The final prediction, denoted as $y$, is a binary classification result derived from the ensemble model. The process is formalized as follows:
@@
Furthermore, for safety-critical tasks with severe class imbalance like depression detection, ensuring the model’s output probabilities are reliable is crucial. The raw scores from a GBT model may be miscalibrated. To address this, a post-processing calibration step can be applied to improve clinical utility. Methods like Isotonic Regression can adjust the classifier’s scores to better reflect true posterior probabilities. This helps correct for potential biases, such as underestimating the probability of the minority class, thereby providing clinicians with more trustworthy confidence scores.
2.6 Design for Explainability
Explainability is critical for safety-sensitive tasks like depression detection (Zhang et al., 2022). While Large Language Models (LLMs) show potential for generating interpretable analysis (Yang et al., 2023a, b), their direct use for classification can be unreliable and sensitive to prompting (Hua et al., 2024). Our hybrid approach is explicitly designed to overcome this by combining a robust traditional classifier with the interpretive power of an LLM, leveraging the strengths of both.
Our system’s explainability is two-fold. First, the final prediction is based on two clinically-relevant, human-readable features generated by an LLM: 1) a set of annotated symptoms aligned with DSM-5 criteria, and 2) a narrative summary of the user’s longitudinal “mood course”. These features serve as direct evidence for the model’s reasoning. Second, to synthesize this evidence for the end-user, the system uses an LLM to generate a final explanatory text, $T^{Exp}$, that explicitly connects the identified symptoms and mood patterns to the classification result ($y$). The prompt used for this generation is detailed in Appendix F.
3 Experimental Setup
3.1 Implementation Details